name two integers whose product is -18 and whose difference between the greatest and least is 9
Answer by oberobic(2304)

x and y are two integers.
.
x*y = -18
.
x-y = 9
x = y+9
.
(y+9)*y = -18
y^2 +9y + 18 = 0
.
(y+3)(y+6) = 0
.
y = -3 or -6, which means x = 6 or 3, respectively
.
The ordered pairs would be: (6,-3) or (3,-6)
.
Done.